# On the convergence of bootstrap current to the Shaing-Callen limit in stellarators

### v1

@@ 4.3 Off-set of bootstrap / Ware pinch coefficient in case of (nearly) aligned maxima
In the advanced stellarator configurations aiming at good confinement of fusion alpha particles, global magnetic field maxima tend to be aligned. This is the case in ideal quasi-symmetric configurations (Nührenberg and Zille, 1988) and also in quasi-isodynamic configurations with poloidally closed contours $B=\rm{const}$ (Mikhailov et al., 2002; Subbotin et al., 2006). For estimations in the latter case, we use a simple model field (94) adding a small perturbation slightly destroying the alignment. In the absence of a perturbation, global maxima are located at $\varphi=2\pi k/N_{\text{tor}}$, with $10<k<N_{\text{tor}}-$ 1, and it is convenient to describe various quantities constant in the off-set wells (toroidal field ripples) such as aspect ratio $A_{j}$, distribution function off-set $g_{\text{off}}^{(j)}$ and weighting factors $w_{\text{off}}^{(j)}$ as functions of the field line label $\vartheta_{0}$ only, omitting the off-set well index $j$. Fixing in stellarator-symmetric fields preserving periodicity $\vartheta_{0}=\vartheta-\iota\varphi+\iota_{p}$ where $\iota_{p}=\pi\iota/N_{\text{tor}}$, we obtain that $g_{\text{off}}^{(j)}=g_{\text{off}}(\vartheta_{0})$ is an odd function which can be cast with help of Eq. (108) for small $\Delta A_{o}\ll 1$ and small mis-alignment of local field maxima $\Delta\eta\ll\delta\eta_{\text{ref}}$ to
@@
$+$ $\left.\Delta^{B}_{j}\left(\Delta\eta\left(\vartheta_{0}-\iota_{p}(2k-1)\right)-\Delta\eta\left(\vartheta_{0}+\iota_{p}(2k-1)\right)\right)\right),$
where both, $\Delta A_{o}(\vartheta_{0})=A_{o}(\vartheta_{0})-1$ and $\Delta\eta(\vartheta)=-\Delta B(\vartheta)/B_{\rm max}^{2}$ are even functions, and $\Delta B(\vartheta)=B(\vartheta,0)-B_{\rm max}\ll B$ is the perturbation of field maxima. The off-set of the Ware pinch coefficient (109) takes in the limit of infinite (irrational) field-line length the form
@@
where an odd function $I_{\cal V}(\vartheta_{0})$ is given by the second Eq. (114) with $\varphi_{j}=0$ and $\varphi_{j+1}=2\pi/N_{\text{tor}}$, and we used (an odd) Fourier series expansion
@@
For the estimates, we use Eqs. (94) and (95) with $\varepsilon_{0}=\varepsilon_{M}\ll 1$, $\varepsilon_{1}=\varepsilon_{t}\ll\varepsilon_{M}$ and $\varepsilon_{3}=0$. In the leading order over these amplitudes, only $A_{o}^{1}\approx\varepsilon_{t}/(2\varepsilon_{M})$ harmonic and $\Delta B_{1}$ harmonic contribute to $\lambda_{\text{off}}$ because of the dominance in the $I_{\cal V}$ spectrum of harmonic $I_{\cal V}^{1}\approx-8\varepsilon_{t}(2\varepsilon_{M})^{1/2}(B_{0}^{2}N_{\text{tor}})^{-1}$. Estimating $l_{c}=\pi R/(2\nu_{\ast})$, $\langle B^{2}\rangle\approx B_{0}^{2}$, $\eta_{b}\approx 1/B_{0}$, $\langle|\lambda|\rangle_{\eta=\eta_{b}}\approx\pi^{-1}(8\varepsilon_{M})^{1/2}$ and $\delta\eta_{\rm ref}\sim 4\pi^{-1/2}(2\varepsilon_{M})^{1/4}(\nu_{\ast}/N_{\text{tor}})^{1/2}B_{0}^{-1}$ we get
@@
It should be reminded here that the off-set (134) due to violation of ripple equivalence appears only for relatively strong alignment of maxima, $\Delta B/B_{0}\lesssim(\nu_{\ast}/N_{tor})^{1/2}\varepsilon_{M}^{1/4}$, and the validity of estimate (134) requires the respective strong inequality. In the marginal case where an approximate equality is realized, the largest off-set, $\lambda_{\text{off}}\sim\lambda_{bB}^{\text{tok}}\;\varepsilon_{t}^{1/2}\varepsilon_{M}^{1/4}\left(\nu_{\ast}N_{\text{tor}}\right)^{-1/2}$ occurs. The result of small mis-alignment of maxima obtained by adding to (94) in case 3 a small perturbation $\Delta B_{1}\cos(\vartheta)$ is shown in Fig. 26. Note that validity condition of linear asymptotic (134) becomes violated for finite $\Delta B_{1}$ at $\nu_{\ast}=1.7\cdot 10^{-8}$ where $\Delta\eta=\delta\eta_{\text{ref}}$, and the case of nearly aligned maxima starts a transition to the usual case of a global maximum reached at a point.
@@ 4.4 Bootstrap resonances
As one can conclude from Fig. 18, resonance structure of $\iota$ (equivalently, of radial) dependence of bootstrap coefficient is absent in finite collisionality cases. This, however, does not mean that these resonances never play a role. The reason for them is the following (see also the Appendix of (Boozer and Gardner, 1990)). By our assumption, magnetic configuration has embedded flux surfaces everywhere so that magnetic islands are absent at any rational magnetic surface, what is manifested by a true surface condition (40). Although these configurations constitute a set of measure zero, they are physically possible. It is straightforward to check that the first term (representing the Pfirsch-Schlüter current) in Eq. (42) is never resonant as long as conditions (40) are fulfilled. However, this is not the case for the second term. Up to a flux surface function, the sub-integrand of the field line integral there is the derivative of the source term $s_{(1)}$, Eq. (3), which fulfills the quasi-Liouville’s theorem (15) at our representative field line, but not necessarily on the other field lines from the same rational surface. In other words, the amplitudes of resonant harmonics fulfilling $\iota m+n=0$ for our rational surface are generally finite, but these are harmonics of $\sin(m\vartheta+n\varphi)=\sin(m\vartheta_{0}+n\varphi_{0}+(\iota m+n)(\varphi-\varphi_{0}))$ which are identical zeros for the starting point $(\vartheta_{0},\varphi_{0})$ being a stellarator symmetry point. Would those amplitudes be all zero, the integral over any field line would be a single-valued function fulfilling then a quasi-Liouville’s theorem (15) for passing particles and, as a consequence, conditions $\oint{\rm d}lv_{\parallel}=\text{const}$ evaluated for passing particles along closed field lines (see the discussion by Shaing and Callen (1983) after Eq.(44b)), which actually mean “true surface” conditions for the effective magnetic field ${\bf B}^{\ast}$ (Morozov and Solov’ev, 1966), i.e. for drift surfaces. The latter, however, need not have the same topology as the magnetic surfaces (see, e.g., (Heyn et al., 2012)) and, at best, can be made embedded for one particular $\eta$ value for the cost of destroyed magnetic surfaces and drift surfaces with different $\eta$ (actually, true drift surface conditions reduce for $\eta=0$ to true magnetic surface conditions). Thus, drift islands are generally present at any rational drift surface (see, e.g., (Smirnova, 1997)), and, due to the increased spectral width of $s_{(1)}$ near the trapped-passing boundary (Boozer and Gardner, 1990) where the exponential decay of $s_{(1)}$ spectrum is replaced with a power law, $(mB_{\varphi}-nB_{\vartheta})(m^{2}+\beta n^{2})^{-3/2}$, drift islands overlap near this boundary ($s_{(1)}=-{\rm d}r/{\rm d}\varphi$ along the orbits by definition (4)). This leads to barely trapped alpha particle losses observed in realistic stellarator configurations (Albert et al., 2020a, b, 2023), which do not ideally fulfil the quasi-symmetry conditions.
@@
Due to our choice of reference field lines, respective drift orbits correspond to invariant axes (X or O points of drift islands), i.e. exact resonance does not destroy their closure and, therefore, dos not affect collisionless $\lambda_{bB}$. However, neighboring resonances result in near resonant contributions to the collisionless $\lambda_{bB}\propto(\iota m-n)^{-1}$ where the denominator is rather close to zero, $\min(\iota m-n)=m\Delta\iota\lesssim 1/n$. One should note that exponential decay of the spectrum is restored from the power law decay for passing particles displaced from the separatrix by $\Delta\eta=\eta-\eta_{b}$ and mode numbers fulfilling $m^{2}+\beta n^{2}>(\eta_{b}^{2}/\Delta\eta)\partial^{2}B/\partial\vartheta^{2}$. Therefore, contribution of near-resonant modes with high $m$ and $n$ to the integral over $\eta$ in Eq. (42) scales with a maximum $\Delta\eta$ which still allows the power law decay, i.e. it is $\propto m^{-1}\Delta\iota^{-1}(m^{2}+\beta n^{2})^{-2}(mB_{\varphi}-nB_{\vartheta})\propto\Delta\iota^{-1}m^{-4}$. Would the rational numbers be evenly distributed over the real axis as we assumed for simplicity in Section 4.2, then $\Delta\iota\sim 1/m^{2}$, and the contribution of near-resonant modes is $\propto m^{-2}$, i.e. it is relatively small. However, uniform distribution is not the case in reality (see, e.g., Fig. 5 in (Kasilov et al., 2002)). Namely, if $\lambda_{bB}$ asymptotic is evaluated at a high-order rational surface with $\iota m_{h}+n_{h}=0$, the distance to the low order resonance $\iota m_{l}+n_{l}=0$ can be as small as $\Delta\iota\sim 1/m_{h}^{2}$ which results in contribution scaling as $m_{h}^{2}/m_{l}^{4}$, i.e. it tends to infinity if the evaluation is at the irrational surface with $\iota$ arbitrarily close to $-n_{l}/m_{l}$. As a result, $\lambda_{bB}$ asymptotic has a fractal dependence on $\iota$, as seen from the right plot in Fig. 18. On the contrary, if $\lambda_{bB}$ asymptotic is evaluated at low order rational field lines, there are no other low order rationals in the neighborhood, and the resulting data points appear to belong to the smooth curve (see the respective markers in the left plot of Fig. 18 which correspond to rational $\iota$ with numerators not larger than 500 while the rest data points have numerators up to 10000).

### v3

@@ 4.3 Off-set of bootstrap / Ware pinch coefficient in case of (nearly) aligned maxima
In the advanced stellarator configurations aiming at good confinement of fusion alpha particles, global magnetic field maxima tend to be aligned. This is the case in ideal quasi-symmetric configurations (Nührenberg and Zille, 1988) and also in quasi-isodynamic configurations with poloidally closed contours $B=\rm{const}$ (Mikhailov et al., 2002; Subbotin et al., 2006). If these configurations are not realized exactly, poloidal (helical in the general case of quasi-symmetry) closure of $B=\rm{const.}$ contours is destroyed, in particular, near the field maximum where these contours split into islands centered around the global maxima reached at one or few points instead of lines. Local maxima ${\bf h}\cdot\nabla B=0$, nevertheless, stay on poloidally closed contours defined in periodic Boozer coordinates by $\varphi=\varphi_{k}^{\text{loc}}(\vartheta)$ where $\varphi_{k}^{\text{loc}}(\vartheta)=\varphi_{0}^{\text{loc}}(\vartheta)+2\pi k/N_{\text{tor}}$ with $k$ being an integer, and $\varphi_{0}^{\text{loc}}(\vartheta)$ is periodic. For simplicity, we restrict our analysis here to the case of poloidally closed contours where also the family of anti-sigma configurations (90) belongs as a sub-class with $\varphi_{0}^{\text{loc}}(\vartheta)=0$. We assume that these local maxima are close to global such that perturbation $\Delta B_{\text{loc}}(\vartheta)\equiv B(\vartheta,\varphi_{0}^{\text{loc}}(\vartheta))-\bar{B}_{\text{max}}$ of the global maximum $\bar{B}_{\text{max}}$ achieved in the ideal (unperturbed) configuration on a line is small. Since $(\vartheta,\varphi)=(0,0)$ is a stellarator symmetry point associated with field maximum, function $\varphi_{0}^{\text{loc}}(\vartheta)$ is odd and $\Delta B(\vartheta)$ is even.
It is convenient now to label the field lines by the poloidal angle in the middle of the first field period where $\varphi=\varphi_{s}\equiv\pi/N_{\text{per}}$ such that $(\vartheta,\varphi)=(0,\varphi_{s})$ is a stellarator symmetry point associated with field minimum of the ideal configuration. Namely, we label them with $\vartheta_{0}=\vartheta-\iota(\varphi-\varphi_{s})=\vartheta-\iota\varphi+\iota_{p}$ where $\iota_{p}=\iota\varphi_{s}$. Positions of local maxima on the so labelled field line, $\varphi_{k}=\varphi_{k}(\vartheta_{0})$, being the solutions to $\varphi_{k}=\varphi_{k}^{\text{loc}}\left(\vartheta_{0}+\iota(\varphi_{k}-\varphi_{s})\right)$ for $-\infty<k<\infty$ can all be expressed via the left maximum of the first period $\varphi_{0}$ on the re-labelled field line as $\varphi_{k}(\vartheta_{0})=\varphi_{0}\left(\vartheta_{0}+2k\iota_{p}\right)+2k\varphi_{s}$. Respective perturbations $\Delta B_{(k)}(\vartheta_{0})$ of global maximum at these points can be expressed via such perturbation $\Delta B_{(0)}$ of the left maximum in the first period as $\Delta B_{(k)}(\vartheta_{0})=\Delta B_{(0)}(\vartheta_{0}+2k\iota_{p})$. Finally, integrals (26) where $\varphi_{j}^{-}=\varphi_{j}$ and $\varphi_{j}^{+}=\varphi_{j+1}$ can be mapped to the first period similarly, $I_{k}(\vartheta_{0},\eta)=I_{0}(\vartheta_{0}+2k\iota_{p},\eta)$.
Due to stellarator symmetry, function $\varphi_{0}(\vartheta_{0})$ is odd with respect to $\vartheta_{0}-\iota_{p}$, i.e. $\varphi_{0}(2\iota_{p}-\vartheta_{0})=-\varphi_{0}(\vartheta_{0})$, while function $\Delta B_{(0)}$ is even. It convenient to express the latter via an even function of the argument $\Delta B(\vartheta^{\prime})$ as $\Delta B_{(0)}(\vartheta_{0})=\Delta B(\vartheta_{0}-\iota_{p})$, such that for anti-sigma configurations we have an identity $\Delta B(\vartheta)=\Delta B_{\text{loc}}(\vartheta)=B(\vartheta,0)-\bar{B}_{\text{max}}$. We fix $\bar{B}_{\text{max}}$ by the condition of zero average, $\overline{\Delta B}=0$ such that an even function $\Delta\eta(\vartheta)=-\Delta B(\vartheta)/\bar{B}_{\rm max}^{2}$ has zero average too. Using definition (84) of the aspect ratio, its perturbation is given by $\Delta A_{o}(\vartheta_{0})=\left(I_{\text{ref}}/I_{0}(\vartheta_{0},\eta_{b})\right)^{1/2}-1$ being an even function with $I_{\text{ref}}$ determined by the condition of zero average $\overline{\Delta A_{o}}$.
Mapping to the first period of the distribution function off-set given by Eq. (105) for small $\Delta A_{o}\ll 1$ and $\Delta\eta\ll\delta\eta_{\text{ref}}$ is similar to that of the other quantities, $g_{\text{off}}^{(0)}(\vartheta_{0})=g_{\text{off}}^{(k)}(\vartheta_{0}-2k\iota_{p})$. Since Eq. (105) gives the off-set in open-ended system with sources located in a single period, solution for the closed (periodic) system with sources in each period is an odd function given by the sum over all periods of the open ended system,
$g_{\text{off}}(\vartheta)=\sum\limits_{k=-\infty}^{\infty}g_{\text{off}}^{(k)}(\vartheta_{0}-2k\iota_{p}),$
which is expressed via even functions $\Delta A_{o}$ and $\Delta\eta$ as
@@
$+$ $\left.\Delta^{B}_{k}\left(\Delta\eta\left(\vartheta_{0}-\iota_{p}(2k-1)\right)-\Delta\eta\left(\vartheta_{0}+\iota_{p}(2k-1)\right)\right)\right).$
Respectively, the off-set of the Ware pinch coefficient (106) is expressed in the limit of infinite (irrational) field-line length via $g_{\text{off}}$ as follows,
@@
where an odd function $I_{\cal V}(\vartheta_{0})$ is given by the second Eq. (111) with $\varphi_{(j)}^{-}=\varphi_{0}(\vartheta_{0})$ and $\varphi_{(j)}^{+}=\varphi_{1}(\vartheta_{0})$, and we used (an odd) Fourier series expansion
@@
For the estimates, we apply general expressions (125)–(129) to the anti-sigma configuration (90) which corresponds to the perturbed case 3 with $\varepsilon_{0}=\varepsilon_{M}\ll 1$, $\varepsilon_{1}=\varepsilon_{t}\ll\varepsilon_{M}$ and $\varepsilon_{3}=0$,
$B(\vartheta,\varphi)=B_{0}\left(1+\varepsilon_{M}\cos(N_{\text{tor}}\varphi)+\varepsilon_{t}\cos\vartheta\left(1-\cos(N_{\text{tor}}\varphi)\right)\right)+\Delta B_{1}\cos\vartheta.$ (131)
In the leading order over $\varepsilon_{t}/\varepsilon_{M}$ and $\Delta B_{1}/B_{0}$, only $A_{o}^{1}\approx\varepsilon_{t}/(2\varepsilon_{M})$ harmonic and $\Delta B_{1}$ harmonic contribute to $\lambda_{\text{off}}$ because of the dominance in the $I_{\cal V}$ spectrum of harmonic $I_{\cal V}^{1}\approx-8\varepsilon_{t}(2\varepsilon_{M})^{1/2}(B_{0}^{2}N_{\text{tor}})^{-1}$. Estimating $l_{c}=\pi R/(2\nu_{\ast})$, $\langle B^{2}\rangle\approx B_{0}^{2}$, $\eta_{b}\approx 1/B_{0}$, $\langle|\lambda|\rangle_{\eta=\eta_{b}}\approx\pi^{-1}(8\varepsilon_{M})^{1/2}$ and $\delta\eta_{\rm ref}\sim 4\pi^{-1/2}(2\varepsilon_{M})^{1/4}(\nu_{\ast}/N_{\text{tor}})^{1/2}B_{0}^{-1}$ we get
@@
It should be reminded here that the off-set (132) due to violation of ripple equivalence appears only for relatively strong alignment of maxima, $\Delta B/B_{0}\lesssim(\nu_{\ast}/N_{tor})^{1/2}\varepsilon_{M}^{1/4}$, and the validity of estimate (132) requires the respective strong inequality. In the marginal case where an approximate equality is realized, the largest off-set, $\lambda_{\text{off}}\sim\lambda_{bB}^{\text{tok}}\;\varepsilon_{t}^{1/2}\varepsilon_{M}^{1/4}\left(\nu_{\ast}N_{\text{tor}}\right)^{-1/2}$ occurs. The result of a small mis-alignment of maxima due to a finite perturbation $\Delta B_{1}$ in Eq. (131) is shown in Fig. 30. Note that the validity condition of a linear asymptotic (132) becomes violated for finite $\Delta B_{1}$ at $\nu_{\ast}=1.7\cdot 10^{-8}$ where $\Delta\eta=\delta\eta_{\text{ref}}$, and the case of nearly aligned maxima starts a transition to the usual case of a global maximum reached at a point.
@@ 4.4 Bootstrap resonances
As one can conclude from Fig. 20, resonance structure of $\iota$ (equivalently, of radial) dependence of bootstrap coefficient is absent in finite collisionality cases (see also Fig. 16 of (Landreman et al., 2022)). This, however, does not mean that these resonances never play a role. The reason for them is the following (see also the Appendix of (Boozer and Gardner, 1990)). By our assumption, magnetic configuration has embedded flux surfaces everywhere so that magnetic islands are absent at any rational magnetic surface, what is manifested by a true surface condition (40). Although these configurations constitute a set of measure zero, they are physically possible. It is straightforward to check that the first term (representing the Pfirsch-Schlüter current) in Eq. (42) is never resonant as long as conditions (40) are fulfilled. However, this is not the case for the second term. Up to a flux surface function, the sub-integrand of the field line integral there is the derivative over $\eta$ of the source term $s_{(1)}$, Eq. (3), which fulfills the quasi-Liouville’s theorem (15) at our representative field line, but not necessarily on the other field lines from the same rational surface. In other words, the amplitudes of resonant harmonics fulfilling $\iota m+n=0$ for our rational surface are generally finite, but these are harmonics of $\sin(m\vartheta+n\varphi)=\sin(m\vartheta_{0}+n\varphi_{0}+(\iota m+n)(\varphi-\varphi_{0}))$ which are identical zeros for the starting point $(\vartheta_{0},\varphi_{0})$ being a stellarator symmetry point. Would those amplitudes be all zero, the integral over any field line would be a single-valued function fulfilling then a quasi-Liouville’s theorem (15) for passing particles and, as a consequence, conditions $\oint{\rm d}lv_{\parallel}=\text{const}$ evaluated for passing particles along closed field lines (see the discussion by Shaing and Callen (1983) after Eq.(44b)), which actually mean “true surface” conditions for the effective magnetic field ${\bf B}^{\ast}$ (Morozov and Solov’ev, 1966), i.e. for drift surfaces. The latter, however, need not have the same topology as the magnetic surfaces (see, e.g., (Heyn et al., 2012)) and, at best, can be made embedded for one particular $\eta$ value for the cost of destroyed magnetic surfaces and drift surfaces with different $\eta$ (actually, true drift surface conditions reduce for $\eta=0$ to true magnetic surface conditions). Thus, drift islands are generally present at any rational drift surface (see, e.g., (Smirnova, 1997)), and, due to the increased spectral width of $s_{(1)}$ near the trapped-passing boundary (Boozer and Gardner, 1990) where the exponential decay of $s_{(1)}$ spectrum is replaced with a power law, $(mB_{\varphi}-nB_{\vartheta})(m^{2}+\beta n^{2})^{-3/2}$ estimated similarly to Eq. (122), drift islands overlap near this boundary ($s_{(1)}=-{\rm d}r/{\rm d}\varphi$ along the orbits by definition (4)). This leads to barely trapped alpha particle losses observed in realistic stellarator configurations (Albert et al., 2020b, a, 2023; Chambliss et al., 2024), which do not ideally fulfil the quasi-symmetry conditions.
@@
Due to our choice of reference field lines, respective drift orbits correspond to invariant axes (X or O points of drift islands), i.e. exact resonance does not destroy their closure and, therefore, does not affect collisionless $\lambda_{bB}$. However, neighboring resonances result in near resonant contributions to the collisionless $\lambda_{bB}\propto(\iota m-n)^{-1}$ where the denominator is rather close to zero, $\min(\iota m-n)=m\Delta\iota\lesssim 1/n$. One should note that exponential decay of the spectrum is restored from the power law decay for passing particles displaced from the separatrix by $\Delta\eta=\eta-\eta_{b}$ and mode numbers fulfilling $m^{2}+\beta n^{2}>(\eta_{b}^{2}/\Delta\eta)\partial^{2}B/\partial\vartheta^{2}$. Therefore, contribution of near-resonant modes with high $m$ and $n$ to the integral over $\eta$ in Eq. (42) scales with a maximum $\Delta\eta$ which still allows the power law decay, i.e. it is $\propto m^{-1}\Delta\iota^{-1}(m^{2}+\beta n^{2})^{-2}(mB_{\varphi}-nB_{\vartheta})\propto\Delta\iota^{-1}m^{-4}$. Would the rational numbers be evenly distributed over the real axis as we assumed for simplicity in Section 4.2, then $\Delta\iota\sim 1/m^{2}$, and the contribution of near-resonant modes is $\propto m^{-2}$, i.e. it is relatively small. However, uniform distribution is not the case in reality (see, e.g., Fig. 5 in (Kasilov et al., 2002)). Namely, if $\lambda_{bB}$ asymptotic is evaluated at a high-order rational surface with $\iota m_{h}+n_{h}=0$, the distance to the low order resonance $\iota m_{l}+n_{l}=0$ can be as small as $\Delta\iota\sim 1/m_{h}^{2}$ which results in contribution scaling as $m_{h}^{2}/m_{l}^{4}$, i.e. it tends to infinity if the evaluation is at the irrational surface with $\iota$ arbitrarily close to $-n_{l}/m_{l}$. As a result, $\lambda_{bB}$ asymptotic has a fractal dependence on $\iota$, as seen from the right plot in Fig. 20. On the contrary, if $\lambda_{bB}$ asymptotic is evaluated at low order rational field lines, there are no other low order rationals in the neighborhood, and the resulting data points appear to belong to the smooth curve (see the respective markers in the left plot of Fig. 20 which correspond to rational $\iota$ with numerators not larger than 500 while the rest data points have numerators up to 10000).